# The Dark Side of Rich Rewards: Understanding and Mitigating Noise in VLM Rewards

### v4

@@ 7.1 Main Results
In Montezuma, Pixl2R$+$BiMI demonstrated 14% performance increase compared to the original models (see Table 2), which is slightly below our expectations. We attribute this result to BiMI’s intentional strategy of providing less frequent discrete rewards. While this strategy effectively reduces false positives, it does not substantially mitigate the inherent reward sparsity issue in Montezuma. However, we discovered a remarkable synergy between BiMI and intrinsic reward models. While previous models showed no significant improvements with DEIR (the intrinsic reward model) alone, combining BiMI and DEIR led to a 65% performance gain. The gap in collaboration effectiveness can be attributed to two factors. In the previous setup, the consistent presence of false positive rewards misled agents towards unacceptable behaviors and hindered further exploration. Now, BiMI’s less frequent but more meaningful rewards provide anchor points for the agent’s learning. Meanwhile, DEIR’s intrinsic rewards fill the gaps between these anchor points, encouraging the agent to explore efficiently in the interim.
@@
See Figure 6 for a quantitative analysis: BiMI rewards are now concentrated on key locations. A significant improvement is the minimal rewards given for falling off cliffs, which was a common source of false positives in the original model. Figure 7 demonstrates a higher success rate in grabbing the key in the first room, one of the most difficult tasks in Montezuma, highlighting the effectiveness of the proposed reward function and its synergy with intrinsic reward models in guiding agents to solve difficult sparse-reward tasks. See results for other environments in Appendix A.10.
7.2 Overall Performance and Ablation Study
@@
The overall improvements were substantial. As shown in Table 2, BiMI led to a 67% improvement for Pixl2R (Markovian) and a 22% improvement for ELLM- (Non-Markovian). These results are also illustrated in Figure 17 in Appendix A.10. Our ablation study highlights the distinct contributions of the binary reward (Bi) and Mutual Information (MI) components within the BiMI framework. The binary reward mechanism alone accounted for a substantial 36.5% improvement in performance. When excluding the results from Crafter, MI component further contributes a 23% improvement over the binary reward alone.
8 Conclusion
We reveal two findings in VLM-based reward functions for RL agents: (1) false positive rewards, rather than false negatives, are more detrimental to policy learning; and (2) our proposed BiMI reward function advocates for pessimistic rewarding, significantly mitigating the slowdown in learning caused by false positives. Our results are supported by both theoretical analysis and empirical validation across three challenging embodied tasks.
9 Limitations
@@
Our study primarily focused on linear sequences of language instructions, excluding more complex cases. Future research should investigate conditional and ambiguous instructions, which likely introduce additional challenges for VLM-based reward models.
We also did not explore finetuning the VLM during agent training, a useful strategy as discussed by Fu et al. (2024). However, we believe that finetuning large VLMs simultaneously during reinforcement learning (RL) is computationally expensive and may not be practical in real-world scenarios.
We acknowledge that mutual information (MI) has been previously explored as a reward objective in RL. However, our work uncovers a novel insight: MI is particularly well-suited for preventing over-reliance on excessively frequent reward sources, which in turn helps to mitigate the issue of too many false positive rewards. This research is innovative in explicitly addressing the practical challenges posed by false positive rewards in instruction-following learning agents that utilize VLM rewards. We believe that our proposed BiMI reward function represents a novel application of MI specifically tailored for VLM-based rewards.
There is a gap in providing a rigorous theoretical foundation for why our theoretical findings extend to non-Markovian reward models like ELLM-. Our Assumption 4.3 facilitates the integration of non-Markovian reward models into the HuRL framework, though some might argue this assumption is strong. However, with advancements in deep RL, the distinction between non-Markovian and Markovian models has become increasingly blurred. For instance, the pioneering DQN (Mnih, 2013) utilized sequences of past observations as input, ostensibly making it non-Markovian. Yet, with deep learning’s capacity to encapsulate complex state histories into a representation that can be treated as Markovian at a higher abstraction level, some have argued that non-Markovian elements can be effectively reconsidered as Markovian within this new context (Hausknecht and Stone, 2015).
Our experiments were conducted in extremely sparse reward settings, with tasks intentionally designed to be challenging by requiring long-horizon planning and sequential execution. This setup may overemphasize the impact of false positive rewards. We acknowledge that in certain cases, rewarding partially correct trajectories may be beneficial for exploration. However, to prevent the agent from overvaluing these partial solutions, it is essential that the reward signal remains conservative, providing a pessimistic estimate of the value of partially correct trajectories.
Appendix A Technical Appendix
Continued from the main text of Understanding and Mitigating Noise in VLM Rewards, the technical appendix consists of the following:
  • §A.1 Detailed Problem Setting, which provides a detailed problem setting of the instruction-following reinforcement learning. This is referred to in Section 3 and Section 4.
  • §A.2 Theoretical Analysis and Proofs, which is referred to in Section 4
  • §A.3 Additional Notes on Theoretical Setting
  • §A.4 Implementation Details of the Experiments, which specify the implementation details for both the first and second stage experiments in Section 5.1 and Section 7 respectively.
  • §A.5 Additional Details of the Experiments of False Positive Rewards, which is referred in Section 5.1.
@@ Appendix A Technical Appendix
  • §A.6 Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores, which provides extra figures that cannot be in the main body due to page limit.
  • §A.7 Impact of Reward Model Quality on Policy Learning Efficiency: Supplementary Experiments and Quantitative Analysis in Montezuma, which provides extra evaluation results to further illustrate false positive rewards hinders agent learning.
  • §A.8 Pseudo-code for Empirical Quantile Calculation for Binary Signal Threshold, which is referred in Section 6.1.
  • §A.9 Additional Implementation Details of the Experiments of BiMI Reward Function, which is referred in Section 7.
  • §A.10 Detailed Experiment Results of BiMI Reward Function, which provides detailed experiment results of BiMI reward function, which is referred in Section 7.
  • §A.11 Deprecated Convergence Analysis, It is retained here solely for reference and backtracking purposes. As the more rigorous theoretical analysis is constructed, the content of this section is now deprecated.
A.1 Detailed Problem Setting
A.1.1 The Pointer Mechanism in Reward Machine
In our problem setting, $r^{v}(\tau_{t},l_{m(t)})$ is the VLM based reward at time step $t$ evaluated using the sub-trajectory $\tau_{t}$ is the sub-trajectory up to time $t$.
@@ A.1.1 The Pointer Mechanism in Reward Machine
Existing VLM-reward implementations applied a pointer mechanism that decide which instruction should the agent consider at the current step. We follow this implementation and denote the pointer as $m(t)$. It indicates the current instruction of the agent is trying to complete at time step $t$. $m(t)$ is updated according to the following rule:
@@
In both the ELLM and the Pixl2R framework that uses VLM-based rewards, the assessment of whether instruction $l_{m(t)}$ is completed at time $t$ is based on the cumulative reward for that instruction. The pointer remains on the current instruction until the accumulated reward reaches a predetermined threshold. For a detailed pseudo-code of the existing VLM reward + RL algorithm, please refer to Appendix A.4.2. Note that our proposed BiMI reward function uses a different mechanism to determine if instruction $l_{m(t)}$ is completed at time $t$.
In logic, an atomic sentence is a type of statement which cannot be broken down into other simpler sentences. While the pointer mechanism effectively enforces the sequential ordering of instructions at a higher level, each individual instruction within the expert walkthrough is not necessarily atomic. This means that a single instruction can encapsulate multiple finer-grained requirements for the agent’s behavior. Therefore, even though the pointer mechanism is implemented, the internal composition of each instruction is not strictly enforced. Consequently, the issue of “composition insensitivity” can arise when the VLM tries to align agents’ trajectories with non-atomic instruction sentences.
Just as potential-based reward shaping (PBRS) modifies the original MDP into a reshaped MDP through the use of auxiliary rewards, the HuRL framework similarly transforms the original MDP $\mathcal{M}$ into a new MDP ~$\widetilde{\mathcal{M}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},\widetilde{r},\widetilde{\gamma}\rangle$, employing a heuristic-based auxiliary reward $h$, along with a coefficient $\beta\in[0,1]$ that scales the auxiliary rewards. Thus, the updated reward function can be written as:
$\widetilde{r}(s,a)=r^{e}(s,a,s^{\prime})+(1-\beta)\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a}[h(s^{\prime})]\quad\text{and $\widetilde{\gamma}=\beta\gamma$}$
We also provide the formal definition of false negative reward as below.
Definition A.1 (False Negative Rewards).
A false negative reward occurs when:
Instruction-Following Perspective: The VLM-based reward $r^{v}(\tau_{t},l_{m(t)})\approx 0$ for a trajectory $\tau_{t}$ that does satisfy instruction $l_{m(t)}$.
A.1.2 Detailed Terminology
This section contains detailed terminology from both the standard RL and the HuRL framework.
We start by setting clear the terminology. The original MDP is defined by a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},r^{e},\gamma\rangle$ as set out in the main text. Let $G^{e}(\tau)$ denote the cumulative environment reward of a single trajectory $\tau=\langle s_{0},a_{0},\ldots,s_{T}\rangle$ where $T$ is the total time of the trajectory, we have
$G^{e}(\tau)=\sum_{t=0}^{T}\gamma^{t}r^{e}(s_{t},a_{t}).$ (3)
A policy $\pi_{\theta}$ is a conditional distribution $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ parameterized by $\theta$, where $\Delta(\cdot)$ denotes the space of probability distribution. The probability of a trajectory $\tau$ being generated, given that actions come from $\pi_{\theta}$ and the starting state is $s_{0}$, is expressed as:
$P(\tau\mid\pi_{\theta})=\prod_{t=0}^{T}\mathcal{P}(s_{t+1}\mid s_{t},a_{t})\pi_{\theta}(a_{t}\mid s_{t})$ (4)
Because a policy $\pi$ is stochastic, different trajectories could be generated from the same policy for a given initial state. The objective function to be optimized is the expected cumulative environment reward of trajectories generated by $\pi$ starting from state $s_{0}$. It is also called the state value function of $s_{0}$ and denoted as:
$V^{\pi}(s_{0}):=\mathbb{E}_{\rho_{s}^{\pi}}\left[G^{e}(\tau)\right]=\sum_{\textnormal{all $\tau_{i}$ generated by $\pi_{\theta}$}}G^{e}(\tau_{i})\cdot P(\tau_{i}|\pi)$ (5)
where $\rho_{s}^{\pi}$ denotes the trajectory distribution of $s_{0},a_{0},s_{1},...$ induced by running $\pi$ starting from $s_{0}=s$. We defined an acceptable policy $\pi_{A}$ as
$\pi_{A}:V^{\pi_{A}}\geq V^{\pi^{*}}-\epsilon$ (6)
where $\pi^{*}$ is the optimal policy and $\epsilon$ is a control parameter that quantifies the acceptance of other policies. Acceptable policies defined in this way are also referred to as $\epsilon$-optimal policies. We set $V^{*}=V^{\pi^{*}}$ for easy reference. Any policy that is not an acceptable policy is an unacceptable policy $\pi_{U}$. We define a good trajectory $\tau_{G}$ as
$\tau_{G}:G^{e}(\tau_{G})\geq V^{*}-\epsilon.$ (7)
Any trajectory that is not a good trajectory is a bad trajectory $\tau_{B}$. Note that this way of distinguishing good/bad trajectories and acceptable/unacceptable policies is explicitly based on the environment reward $r^{e}$, and introducing additional or alternative rewards into the system does not change this definition. For easy reference, we call any trajectory that reaches a goal state a goal trajectory, and we call any policy that can generate at least one goal trajectory a goal policy.
We denote the state distribution of a policy $\pi$ at time $t$ as $d^{\pi}_{t}$. Thus, the discounted average state distribution of a policy $\pi$ can be expressed as $d^{\pi}=(1-\gamma)\sum^{\infty}_{t=0}\gamma^{t}d^{\pi}_{t}$, where $(1-\gamma)$ is the normalization factor to ensure the result is a proper probability distribution. We also define a shorthand notation – for a state distribution $d\in\Delta(\mathcal{S})$, we define $V(d)=\mathbb{E}_{s\sim d}\left[V(s)\right]$.
When the Heuristic-Guided Reinforcement Learning (HuRL) framework is applied, the reshaped MDP $\widetilde{\mathcal{M}}$ has separated reshaped reward $\widetilde{r}$ and separated discount factor $\widetilde{\gamma}=\beta\gamma$, where $\beta\in[0,1]$ is the hyperparameter that scales the auxiliary heuristic rewards. We denote the value function of the reshaped MDP as $\widetilde{V}$, and the optimal policy and its value function under $\widetilde{\mathcal{M}}$ as $\widetilde{\pi}^{*}$ and $\widetilde{V}^{*}(\widetilde{V}^{\widetilde{\pi}^{*}})$.
The discussion regarding the convergence guarantee of learning $\pi$ in $\widetilde{\mathcal{M}}$ resulting in the optimal policy in the original MDP $\mathcal{M}$ can be found in Appendix A.3.1. For the purposes of this paper, it is sufficient to note that the authors of HuRL have provided a trick to ensure convergence. Moreover, empirical experiments conducted by the authors demonstrate that HuRL converges even without this trick.
It is also important to distinguish the Bellman backup equation under the two different MDPs: the original $\mathcal{M}$ and the reshaped $\widetilde{\mathcal{M}}$. By definition, $(\mathcal{B}h)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$. In contrast, (~$(\widetilde{\mathcal{B}}h)(s,a)=\widetilde{r}(s,a)+\widetilde{\gamma}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$. Nevertheless, the two Bellman backup equations possess a remarkable property – they are essentially equivalent to each other:
(~$(\widetilde{\mathcal{B}}h)(s,a)$ $=\widetilde{r}(s,a)+\widetilde{\gamma}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
$=(\mathcal{B}h)(s,a)$ (8)
A.1.3 Detailed Construction of Markovian Heuristic using Non-Markovian Rewards from VLM
Specifically,
$h(s_{t}):=\begin{cases}A(s_{t})V^{*}(s_{t})&\text{if $s_{t}\in S^{m(t)}$}\\ 0&\text{otherwise}\end{cases}$ (9)
where
$A(s_{t}):=\frac{\Pi_{t^{\prime}=0}^{t}r^{v}(\tau_{t^{\prime}},l_{m(t^{\prime})})}{\Pi_{t^{\prime}=0}^{t}r^{v}(\tau^{*}_{t^{\prime}},l_{m(t^{\prime})})}\in[0,1]$ (10)
is a scaling factor that quantifies how closely the agent is following previous instructions.
A.2 Theoretical Analysis and Proofs
A.2.1 Convergence time on a sparse-reward landscape
We can pin down a characteristic convergence time on a sparse-reward landscape. The sparse-reward setting enforces that

### v5

@@ 8 Experiments and Results
Montezuma  Pixl2R++BiMI demonstrated 14% performance increase compared to the original models (see Table 2), which is slightly below our expectations. We attribute this result to BiMI’s intentional strategy of providing less frequent discrete rewards. While this strategy effectively reduces FPs, it does not substantially mitigate the inherent reward sparsity issue in Montezuma. However, we discovered a remarkable synergy between BiMI and intrinsic reward models. While previous models showed no significant improvements with DEIR (the intrinsic reward model) alone, combining BiMI and DEIR led to a 65% performance gain. The gap in collaboration effectiveness can be attributed to two factors. In the previous setup, the consistent presence of false positive rewards misled agents towards unacceptable behaviors and hindered further exploration. Now, BiMI’s less frequent but more meaningful rewards provide anchor points for the agent’s learning. Meanwhile, DEIR’s intrinsic rewards fill the gaps between these anchor points, encouraging the agent to explore efficiently in the interim.
@@
See Figure 6 left and Figure 6 for a quantitative analysis: BiMI rewards are now concentrated on key locations. A significant improvement is the minimal rewards given for falling off cliffs, which was a common source of FPs in the original model. Figure 7 demonstrates a higher success rate in grabbing the key in the first room, one of the most difficult tasks in Montezuma, highlighting the effectiveness of the proposed reward function and its synergy with intrinsic reward models in guiding agents to solve difficult sparse-reward tasks.
Minigrid: ELLM-++BiMI achieved a remarkable 74% improvement in performance compared to the original models.This substantial gain is particularly noteworthy given the unique challenges presented by Minigrid. The abstract, shape-based visuals diverge drastically from the natural images used in VLMs’ pretraining, preventing the models from effectively utilizing their prior pretraining knowledge. Consequently, VLMs struggled to accurately assess similarities between Minigrid’s abstract visuals and textual instructions, resulting in highly noisy reward signals. The significant improvement demonstrated by BiMI underscores its effectiveness in handling noisy signals, directly addressing our primary research challenge. This capability is crucial for deploying instruction-following agents in real-world, unfamiliar scenarios, where visual inputs often deviate from the VLMs’ training distribution, leading to noisy reward signals.
Crafter: We observed an intriguing pattern of results. The Bi component alone led to 14% and 3.2% improvement in performance over the original models. However, contrary to our observations in other environments, the addition of the MI component actually decreased this improvement. This unexpected outcome can be attributed to the unique nature of Crafter task, where agents must repeatedly achieve the same subtasks (e.g., drinking water) for survival. The MI component, designed to discourage over-reliance on frequently occurring signals, inadvertently penalized the necessary repetition of survival-critical actions. Nevertheless, Bi alone still managed to improve performance over vanilla VLM-based reward models, suggesting that reducing FPs is still beneficial across all testing environments. The combination of BiMI with DEIR (the intrinsic reward model) also showed promising results, indicating a productive balance between exploration (driven by DEIR) and exploitation (guided by BiMI instruction reward).
Overall Performance and Ablation Study The BiMI reward function yielded significant performance gains, as detailed in Table 2. For the Markovian Pixl2R model, BiMI improved scores by 67%, while the non-Markovian ELLM- model improved by 22%. These improvements, alongside synergy with intrinsic rewards, are depicted in Figure 16, which illustrates how combining BiMI with intrinsic rewards (i.e., DEIR) boosts agent performance across different environments.
@@
Our ablation study further demonstrates the distinct contributions of the binary reward (Bi) and mutual information (MI) components within the BiMI framework with results shown in Figure 17. The Bi mechanism alone drove a 36.5% performance increase over baseline models. Excluding the Crafter environment, the MI component added a further 23% improvement over Bi alone. Together, these results underscore the importance of both components in mitigating false positive rewards and enhancing agent performance.
9 Conclusion and Discussion
Most existing work on VLM-based reward models for embodied instruction-following agents remains at the proof-of-concept stage in simplified environments (Du et al. 2023; Rocamonde et al. 2024; Wang et al. 2024), overlooking the challenge of noisy rewards in complex, long-horizon tasks. Our work addresses this gap through theoretical analysis grounded in heuristic admissibility — a classical concept from the automated planning domain but rarely applied to embodied learning agents. This perspective enables us to rigorously explain why VLM-based reward models are prone to failure without explicit noise handling. Building on this foundation, we identify two key insights: (1) false positive rewards are substantially more harmful to policy learning than false negatives, and (2) the proposed BiMI reward function, which incorporates pessimistic rewarding, effectively mitigates this problem. Our findings are supported across three challenging embodied tasks, spanning Markovian and non-Markovian reward models.
@@ 9 Conclusion and Discussion
Limitations  We primarily focused on linear sequences of language instructions, excluding more complex cases. Future research should investigate conditional and ambiguous instructions, which likely introduce additional challenges for VLM-based reward models. There is also a gap in providing a rigorous theoretical foundation for why our theoretical findings extend to non-Markovian reward models. However, with advancements in deep RL, the distinction between non-Markovian and Markovian models has become increasingly blurred (Hausknecht and Stone 2015).
Appendix Table of Contents
  1. 1.
    Anticipatory defense on the choice of test environments and metrics (Appendix A)
    Justification for chosen test environments and evaluation metrics.
  2. 2.
    Detailed Related Work (Appendix B)
    An expanded discussion of related research in the domain and research gaps.
  3. 3.
    The Complete Procedure of the VLM-based Reward Model (Appendix C)
    The complete procedure of involving VLM-based reward model in the RL training loop.
  4. 4.
    Convergence time on a sparse-reward landscape (Appendix D)
    Proof of convergence time on a sparse-reward landscape.
  5. 5.
    VLM-based reward model is a heuristic function (Appendix E)
    Proof that VLM-based reward model is a heuristic function.
  6. 6.
    False Positive and the Violation of Pessimistic Property of Heuristics (Appendix F)
    Proof that false positive rewards violate the pessimistic property of heuristics.
  7. 7.
    False Positive gives unbounded bias in the optimality gap (Appendix G)
    Proof that false positive rewards give unbounded bias in the optimality gap.
  8. 8.
    Additional Experimental Setup Details (Appendix H)
    Supplementary information about experimental configurations.
  9. 9.
@@
    Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores (Appendix I)
    Extra figures for stage 1 experiments.
  10. 10.
    Extra Experiment on Simulated Oracle Reward Model to demonstrate H3 and H4 (Appendix J)
    Supporting experiments for hypothesis H3 and H4 using simulated oracle reward models.
  11. 11.
    Algorithm for Empirical Quantile Calculation (Appendix K)
    Algorithm for calculating empirical quantile of cosine similarity scores.
  12. 12.
    Ablation Study Lineplots for BiMI reward (Appendix L)
    Lineplots for ablation study of BiMI reward function.
  13. 13.
    Encoder-based Vision-Language Models Explanation (Appendix M)
    A brief overview of two types of VLMs, encoder-based and generative VLMs.
  14. 14.
    Sparse Reward and Random Walk (Appendix N)
    Proposition and proof of the relationship between RL convergence in sparse reward setting and random walk.
Appendix A Anticipatory defense on the choice of test environments and metrics
As discussed in Appendix B, we need test environments that clearly demonstrate the negative impact of noisy reward signals from VLM-based reward models. This requires complex state spaces (both large and diverse) where the approximation errors of VLMs become evident — specifically, where they struggle to provide accurate alignment scores between an agent’s trajectory and given instructions due to the complexity of visual observations. Additionally, we need longer-horizon tasks where the impact of noisy reward signals can accumulate over time. In contrast, many embodied environments used in previous studies are not suitable due to their short-horizon tasks and simple state spaces. Table 3 lists common RL embodied environments used in the literature, explaining why certain environments are inadequate for our evaluation purposes while highlighting suitable environments we aim to use.
Table 3: Common RL embodied environments and their suitability for evaluating noisy reward signals from VLM-based reward models.
Env Type State Space Complexity Horizon Length Instruction Source
Not Suitable Envs
CartPole (Brockman et al. 2016) Classic Control Simple Short Manual annotation
MountainCar (Brockman et al. 2016) Classic Control Simple Short Manual annotation
Humanoid (Todorov, Erez, and Tassa 2012) Bipedal robot control Complex Short Manual annotation
MetaWorld (Yu et al. 2019) Robotic arm Moderate Short Predefined task descriptions
SoftGym (Lin et al. 2020) Robotic arm Moderate Short Predefined task descriptions
Suitable Envs
Montezuma (Bellemare et al. 2013) Platformer Complex Long Manual annotation
Crafter (Hafner 2022) Open world survival Complex Long Programmable (via engine data)
Minigrid (Chevalier-Boisvert et al. 2023) Grid-world maze Controllable Controllable Predefined task descriptions
We describe each testing environment used in our experiments. More details introduction can be found in on the official project homepage of each benchmark (Hafner 2022; Bellemare et al. 2013; Chevalier-Boisvert et al. 2023).
A.1 Metrics
In our experiments, we used a score metric adapted from the Crafter benchmark (Hafner 2022) to evaluate agent performance across different environments. This score metric aggregates success rates for individual subtasks using a geometric mean. Formally, the score metric is defined as follows:
  1. 1.
  2. 2.
  3. 3.
To evaluate the learning speed of our RL algorithms, we utilize the Area Under the Learning Curve (AUC) metric. This metric, previously employed in the literature (Goyal, Niekum, and Mooney 2019, 2020), quantifies the cumulative performance of the agent throughout the training process, effectively indicating how rapidly the agent improves its policy. We implement a win cap, limiting the maximum number of wins an agent can achieve during training to a certain number (e.g., 5000). Training automatically terminates once this limit is reached, under the assumption that policy learning has stabilized by this point. The AUC is therefore number of wins normalized by the total number of training episodes. Formally, this is expressed as:
$\text{AUC}=\frac{\sum_{i=0}^{T}\text{wins}_{i}}{T}$
where $\text{wins}_{i}$ is the number of wins at episode $i$, and $T$ is the total number of training episodes.
Appendix B Detailed Related Work
The approach of converting natural language instructions into reward signals for RL agents has been a longstanding area of exploration. Early studies (Kaplan et al. 2017; Goyal, Niekum, and Mooney 2019; Wang et al. 2019; Goyal, Niekum, and Mooney 2020) trained separate text and vision encoders from scratch so as to convert multimodal data into continuous feature vectors, then used a discriminator to assess alignment between trajectories and instructions. The cosine similarity metric, found by Frome et al. (2013) as an effective measure of semantic alignment, became foundational to these methods. Strikingly, the core architecture of VLM-based reward models, as illustrated in Figure 1, has remained largely unchanged throughout the years. Both pioneering efforts (Kaplan et al. 2017; Goyal, Niekum, and Mooney 2019; Wang et al. 2019; Goyal, Niekum, and Mooney 2020) and modern implementations continue relying on computing cosine similarity scores between semantic embeddings of instructions and agent trajectories to generate rewards. Formally, given two embedding vectors, $E(l)$ and $E(\tau)$, which denote the semantic embeddings of an instruction $l$ and a trajectory $\tau$, respectively, the cosine similarity is defined as:
$\text{Cosine Similarity}(E(l),E(\tau))=\frac{E(l)\cdot E(\tau)}{\|E(l)\|\|E(\tau)\|}$ (6)
However, before the advent of CLIP (Radford et al. 2021) in 2021, these efforts were hampered by the absence of powerful pre-trained encoder-based VLMs. Consequently, researchers focused on proof-of-concept studies, testing their approaches on simpler, short-horizon tasks or sub-tasks — for example, Goyal, Niekum, and Mooney (2019) evaluated their method on sub-tasks rather than the full Montezuma’s Revenge game333available at https://www.retrogames.cz/play˙124-Atari2600.php. This likely reflected an understanding that complex, long-horizon tasks were impractical at the time. As such, the field had yet to grapple with noisy reward signals, which remained minimal in these simpler environments.
CLIP (Radford et al. 2021), an encoder-based VLM, marked a turning point by providing a high-quality joint semantic embeddings for vision and language data. Leveraging CLIP, researchers started addressing more complex environments. For example, Mahmoudieh, Pathak, and Darrell (2022) used a CLIP-based reward model to guide a robotic arm in manipulation tasks, demonstrating its feasibility for real-world applications.
In addition to the topics mentioned in the main body of the paper, we continue to highlight related work on reward signals and mitigation strategies below.
Reward Signal from Human Preference. Recent work on RL from Human Feedback (RLHF) (Ouyang et al. 2022) also leverage expert preference as a reward signal. However, our work differs in key aspects. Unlike RLHF’s focus on textual outputs, our approach involves evaluating cross-modal similarities between visual and textual data in environments requiring long-horizon decision-making and frequent embodied interactions, a domain not typically covered by RLHF.
Mitigating Misspecified Rewards. Prior works proposed mitigating false positive rewards by training a parallel exploration policy to escape local optima caused by misspecified rewards (Ghosal et al. 2022; Fu et al. 2024). In contrast, we propose a novel reward function that directly penalizes likely false positive reward signals during training. We further show that our method complements exploration-based methods and achieves superior performance when combined.
Research Gap Successes with CLIP-based reward models have been particularly pronounced in environments with short-horizon tasks. For instance, Rocamonde et al. (2024) achieved promising results in classic control tasks such as CartPole, MountainCar (Brockman et al. 2016), and Humanoid (Todorov, Erez, and Tassa 2012), environments that lack long-horizon dependencies. Similarly, Wang et al. (2024) used the SoftGym environment (Lin et al. 2020), which includes a suite of short-horizon robotic arm manipulation tasks, as well as MetaWorld environment (Yu et al. 2019), which lacks strict ordering constraints on subgoals, eventually leading to a lack of long-horizon dependencies.
Importantly, VLM reward models are learned-based reward models and, as such, inherently suffer from noisy reward signals. Sadly, the tendency to focus on simpler environments, or to sidestep the challenges of noisy reward signals, keeps appear in recent work. For example, Chan et al. (2023) explored CLIP-based rewards to reduce dependence on human-engineered reward functions, but their experiments were confined to simple housekeeping tasks with small action and state spaces — far less complex than pixel-based environments like Montezuma’s Revenge. Similarly, ELLM (Du et al. 2023) relied on a hard-coded oracle reward model (labeled as a “VLM reward model” for conceptual demonstration) to maintain training stability. Notably, without this oracle, their agent performed worse than a pure RL agent without any reward shaping — highlighting the limitations imposed by noisy VLM-based rewards in complex environments.
Despite the growing interest of using VLM-based reward models in embodied RL, the literature has largely overlooked the detrimental effects of noisy reward signals on learning efficiency. Key questions remain underexplored: how reliable are these models in complex, long-horizon environments, how do false positive and false negative reward noise affect learning speed, and how to maintain effective learning under noisy reward conditions. These gaps highlight a critical area for progress. In this work, we aim to address this gap by examining these issues in depth and proposing solutions to enhance the applicability of language-guided RL in real-world scenarios.
Appendix C The Complete Procedure of the VLM-based Reward Model
@@
Language instructions in real-world tasks are rarely singular; they typically form a sequence that guides an agent step-by-step. A VLM-based reward model must therefore include a mechanism to transition between instructions as the agent progresses. Existing VLM-reward implementations applied a pointer mechanism to decide which instruction the agent should consider at the current step. We follow this implementation and denote the pointer as $m(t)$. It indicates the instruction that the agent is trying to complete at time step $t$. $m(t)$ is updated according to the following rule:
@@
The pointer remains on the current instruction until the accumulated reward for completing $l_{m(t)}$ reaches a predetermined threshold.
Given a sequence of instruction sentences $L=\{l_{1},l_{2},\dots,l_{n}\}$, a typical VLM-based reward model, as seen in Pixl2R (Goyal, Niekum, and Mooney 2020) and ELLM (Du et al. 2023), maintains this pointer to the current instruction, starting with $l_{1}$. To track progress, it imposes a maximum reward cap — for example, 2.0 — on each instruction. Once the cumulative reward reaches this cap, the pointer advances to the next instruction in the sequence. This approach ensures the agent is incentivized to complete one subtask before moving on to the next.
The complete procedure for training an RL agent with a VLM-based reward model is detailed in algorithm 1. This algorithm builds on concepts including MDP formulation, RL algorithms, and the VLM-based reward model (Appendix B).
Algorithm 1 Instruction-following RL training with VLM-based reward model
1: Initialize policy network $\pi_{\theta}$
2: Initialize value network $Q_{\phi}$
3: Setup VLM-based reward model $E(\cdot)$
4: Split instruction essay into sentences $\{l_{1},l_{2},...,l_{K}\}$
5: Initialize instruction pointer $p=1$
6: Initialize cumulative VLM reward $r_{\text{cum}}=0$
7: Initialize cumulative VLM reward threshold $q$
8: Initialize replay buffer $\mathcal{D}$
9: Initialize agent trajectory memory queue $\tau$ with length $W$
10: for each episode do
11:   Initialize state $s_{0}$
12:   for $t=0$ to T1T-1 do
13:    Select action $a_{t}\sim\pi_{\theta}(a_{t}|s_{t})$
14:    Execute $a_{t}$, observe next state $s_{t+1}$ and environmental reward $r^{e}_{t}$
15:    Enqueue $(s_{t},a_{t},r_{t},s_{t+1})$ in $\tau$
16:    $\triangleright$  Compute VLM reward   $\triangleleft$:
17:    $r^{v}_{t}=p(l_{p}|\tau)=\frac{E(\tau)\cdot E(l_{p})}{\|E(\tau)\|\|E(l_{p})\|}$
18:    Combine rewards: $r_{t}=r^{e}_{t}+(1-\beta)\gamma r^{v}_{t}$ $\triangleright$ $\beta$ is a scaling factor
19:    Store $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{D}$
20:    $r_{\text{cum}}\leftarrow r_{\text{cum}}+r^{v}_{t}$
21:    if $r_{\text{cum}}\geq q$ then
22:      $p\leftarrow\min(p+1,K)$
23:      $r_{\text{cum}}\leftarrow 0$
24:    if Reach Update Frequency then
25:      Sample mini-batch $\{(s_{j},a_{j},r_{j},s_{j+1})\}$ from $\mathcal{D}$
26:      $\triangleright$  Compute TD errors   $\triangleleft$:
27:      $\delta_{j}=r_{j}+\gamma Q_{\phi}(s_{j+1},a_{j+1})-Q_{\phi}(s_{j},a_{j})$
28:      $\triangleright$  Update value network   $\triangleleft$:
29:      $\phi\leftarrow\phi+\alpha_{v}\sum_{j}\delta_{j}\nabla_{\phi}Q_{\phi}(s_{j},a_{j})$
30:      $\triangleright$  Update policy network   $\triangleleft$:
31:      $\theta\leftarrow\theta+\alpha_{p}\sum_{j}Q(s_{j},a_{j})\nabla_{\theta}\log\pi_{\theta}(a_{j}|s_{j})$
Appendix D Convergence time on a sparse-reward landscape
In this work, Section 4, we have shown that the convergence time of a policy learning algorithm on a sparse-reward landscape can be characterized by the distance in the parameter space. Here, we provide a detailed proof of the convergence time on a sparse-reward landscape. We can pin down a characteristic convergence time on a sparse-reward landscape. The sparse-reward setting enforces that